what is:
7/12 minus 6/8 ????

Answer by stanbon(48569)

7/12 minus 6/8
---------------------------
(7/12)-(6/8)
= (7/12)-(3/4)
The least common denominator is 12.
= (7/12)-(9/12)
= -2/12
=-1/6
=============================
Cheers,
Stan H.